OpenStudy (anonymous):

prove cosine(pi*t*t/8) is not periodic

OpenStudy (anonymous):

A function \(f(x)\) is periodic if you can find \(l\) such that \(f(x)=f(x+l)\) with \(l>0\). Assume the given function is periodic, then \[\begin{align*} \cos\left(\frac{\pi t^2}{8}\right)&=\cos\left(\frac{\pi (t+l)^2}{8}\right)\\\\ \frac{\pi t^2}{8}&=\frac{\pi (t+l)^2}{8}\\\\ t^2&=(t+l)^2\\\\ t^2&=t^2+2lt+l^2 \end{align*}\] If \(l>0\), this can't be true. Hence we have a contradiction, so the given function cannot be periodic.

OpenStudy (loser66):

@SithsAndGiggles We have l =0 and l = -2t Those cases show it is periodic. Am I right?

OpenStudy (anonymous):

@Loser66 the period of a function must be constant, so it can't be \(l=-2t\). As for the case \(l=0\), consider the function \(f(x)=x\). This function is periodic if \(f(x)=f(x+l)\), which gives \[x=x+l~~\implies~~l=0\] but we know \(f(x)=x\) is not periodic. As I said in my first post, the period \(l\) must be non-zero.